two dice are rolled find the probability of getting
a. sum of 5,6,7
b.doubles or a sum of 6 or 8
c. a sum greater than 8 or less than 3

Answer by edjones(8007)

Make a box along the outside edge write the numbers 1 through 6 on the top and left edge. now put the 36 sums of the outside numbers inside. Then you will understand the answers below.
.
a) 4/36,5/36,6/36
.
b)6+4+4=14/36 I subtracted 1 each for the 6 and 8 to avoid counting the 3,3 and 4,4 twice.
.
c)4+3+2+1+1=11/36
.
Ed